OpenStudy (anonymous):

Find the derivative of the parametric curve x = t^2 + 2t -1 and y = 3t^3 + 2t^2 - 1 at t = 1

OpenStudy (tkhunny):

\(\dfrac{dy}{dx} = \dfrac{\dfrac{dy}{dt}}{\dfrac{dx}{dt}}\)?

OpenStudy (anonymous):

Thank you!!